OpenStudy (anonymous):

1. What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms? (1 point) 2. What is the sum of the geometric sequence 4, 12, 36 … if there are 9 terms? (1 point) 3. What is the sum of a 6-term geometric sequence if the first term is 11, the last term is –11,264 and the common ratio is –4? (1 point) 4. What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250? (1 point)

OpenStudy (anonymous):

needd help on these 4.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use the geometric series formula

ganeshie8 (ganeshie8):

1. What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms? (1 point)

OpenStudy (anonymous):

sn=a1-a1^rn/1-r . right ?

OpenStudy (mertsj):

That looks arithmetic to me.

OpenStudy (jhannybean):

HAHA YEAH . Mybad.

OpenStudy (jhannybean):

I had a brain fart for a minute.

OpenStudy (mertsj):

All of these questions ask for the sum of a geometric sequence. The formula for that sum is: \[S _{n}=\frac{a _{1}(1-r^n)}{1-r}\]

OpenStudy (mertsj):

For the first one, the first term is 8 so a_1=8 You always multiply by -2 to get the next term so r=-2 There are 15 terms so n = 15. Plug into the formula and simplify.

OpenStudy (anonymous):

what ? i thought geometric sequence was an=a1-a1^rn/1-r

OpenStudy (mertsj):

Ok. Do what you want. I have shown you the formula and how to do the problem but you are perfectly free to do whatever you want.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

forget it .

ganeshie8 (ganeshie8):

yo give it a try